Natalie used 112 inches of blue yarn as a border around her rectangular bulletin board. If the bulletin board is 36 inches wide, how long is it
Answer by checkley77(12844)

P=L*W
112=L*36
L=112/36
L=3.111 INCHES FOR THE LENGTH.
PROOF:
112=3.111*36
112=112